# Asymmetric coupling of non-chaotic Rulkov neurons: fractal attractors, quasi-multistability, and final state sensitivity

### v1

@@ IV.2 Four-dimensional basins
which the slice in Figure 6 only intersects at its center.
@@ V Fractal basin boundary analysis
In Figs. 5 and 6, it is clear that the basins of the non-chaotic spiking attractor and strange pseudo-attractor are not divided by a smooth boundary. This naturally leads us to suspect that the basin boundary $\Sigma$ between the white and black basins might be fractal, which would lead to uncertainty regarding which attractor a given initial state will end up in [28]. In this section, we use the uncertainty exponent method [47] to determine the dimension of the basin boundary $\Sigma$ and the degree to which final states (in the short term) are uncertain. Briefly, the uncertainty exponent method works as follows. Consider a dynamical system in $n$-dimensional state space, and let $\mathbf{x}_{0}$ be an initial state with a small uncertainty $\epsilon$, meaning the true initial state may be anywhere in the $n$-ball $|\mathbf{x}-\mathbf{x}_{0}|<\epsilon$. Let $A$ and $C$ be the only two attractors of the dynamical system, and let $\hat{A}$ and $\hat{C}$ be their associated basins of attraction. $\mathbf{x}_{0}$ will be attracted to either $A$ or $C$, but the uncertainty $\epsilon$ may cause the perturbed initial state to be in a different basin than $\mathbf{x}_{0}$. Then, we expect the probability that we will falsely predict where a random uncertain state in $|\mathbf{x}-\mathbf{x}_{0}|<\epsilon$ will end up, denoted by $\varrho(\epsilon)$, to follow a power law:
@@
where $\mathfrak{u}\leq 1$ is called the uncertainty exponent. Then, when $\mathfrak{u}$ is small, significantly minimizing initial state uncertainty will only have a small effect on final state uncertainty. It is shown in [47] that the uncertainty exponent $\mathfrak{u}$ is related to the dimension of the basin boundary dividing $\hat{A}$ and $\hat{C}$ by
@@
Returning to the asymmetrically coupled Rulkov neuron system, we present the results of our analysis of the basin boundary $\Sigma$ between the white and black basins using uncertainty exponents. Similar to how we classified the intersection of the neuron system’s basins with $S_{2}$ and $S_{4}$, we examine the uncertainty exponents and fractalization of both $\Sigma\cap S_{2}^{\prime}$ and $\Sigma\cap S_{4}^{\prime}$. We have already defined $S_{2}^{\prime}$ as a specific square-shaped subset of $S_{2}$ (Eq. 22), and we will similarly define $S_{4}^{\prime}\subset S_{4}$ to be the four-dimensional hypercube
@@
to see if any of them end up in a different basin than $\mathbf{X}_{0}$ does.
@@
with an $R^{2}$ value of 0.986. This indicates that $\mathfrak{u}_{2}\approx 0.314$, and since this is less than one, final state uncertainty does indeed exist in $S_{2}^{\prime}$. By Eq. 39, the fractal dimension $d_{2}$ of $\Sigma\cap S_{2}^{\prime}$ is
@@
with an $R^{2}$ value of 0.967. This indicates $\mathfrak{u}_{4}\approx 0.037$, which is exceedingly small and indicates extreme final state sensitivity. To put this in perspective, the fact that $\varrho_{4}(\epsilon)$ is proportional to $\epsilon^{\mathfrak{u}_{4}}$ means that to reduce the uncertainty regarding which attractor an initial state in $S_{4}^{\prime}$ will be attracted to by a factor of 10, we will need to reduce the initial uncertainty $\epsilon$ by a factor on the order of $10^{27}$. By Eq. 39, the fractal dimension $d_{4}$ of the basin boundary $\Sigma\cap S_{4}^{\prime}$ is
$d_{4}=n-\mathfrak{u}_{2}\approx 4-0.037=3.963.$ (46)
This indicates that this basin boundary is extremely fractal; it has a comparable geometry to a true four-dimensional object even though it divides four-dimensional space. Even though Fig. 6(b) doesn’t show the entire four-dimensional set $S_{4}^{\prime}$, it justifies this high fractal dimension, where it is clear that the boundary dividing the seemingly random distribution of white and black points must be close to the dimension of the system’s state space.
@@ VI Conclusions
The primary objective of this study was to analyze, classify, and quantify the geometries of the attractors, basins of attraction, and fractal basin boundaries of an asymmetrically electrically coupled system of two identical non-chaotic Rulkov neurons. We discovered the existence of a quasi-multistability in the system, with a true non-chaotic spiking attractor and a chaotic spiking-bursting pseudo-attractor. Using box counting, we found that the chaotic pseudo-attractor was fractal but that its fractal dimension didn’t coincide with its Lyapunov dimension. We also classified the basins of attraction of the non-chaotic spiking attractor and the chaotic pseudo-attractor. We found that, in a two-dimensional slice of the basins near the slow variable values of the attractors, the basin of the chaotic pseudo-attractor took up almost all of the state space slice, while the basin of the non-chaotic spiking attractor had a finite measure and occupied increasingly small fractions of state space far away from its attractor. However, in all of four-dimensional state space, we discovered that both basins occupied fixed fractions of it. Finally, analyzing the system’s uncertainty exponents and basin boundaries, we found extreme final state sensitivity requiring initial state uncertainty improvements on the order of $10^{27}$ to improve final state uncertainty by a factor of 10, as well as fractal basin boundaries with dimension $d\approx 3.96$ taking up almost all of four-dimensional state space.
@@
We believe that our findings of complex and interesting geometries in the basins of attraction of this asymmetrically coupled Rulkov neuron system can be applied to other neuron models and dynamical systems, and we hope to see further study done in the classification of basins of attraction and quantification of fractal basin boundaries and final state sensitivity in the context of continuous-time and discrete-time neuronal systems. We also suggest research be done in using the method of basin entropy [62] to analyze the final state uncertainty and basin geometries of neuron systems, which extends this work and, to the best of our knowledge, has never been done before. Finally, building on existing studies with biological neurons [63, 64, 65], we encourage exploration in experimentally replicating our results regarding final state sensitivity and complex basin geometry with real neurons.
@@ Appendix A Jacobian Matrix Partition
Using this, we can compactify $J(\mathbf{X})$ into the form shown in Eq. 11.

### v3

@@ IV.2 Four-dimensional basins
which the slice in Fig. 7 intersects only at its center.
To get a more comprehensive picture of the four-dimensional basins, we take inspiration from Ref. [61] to present a small three-dimensional slice of the black basin in Fig. 8, where we mark initial states that get attracted to the nonchaotic spiking attractor with a small black point (the empty space represents the white basin). The initial states included in this plot have initial $x_{1},x_{2},y_{1}$ values between $-2\leq x_{1,0}\leq 2$, $-2\leq x_{2,0}\leq 2$, and $-3.5\leq y_{1,0}\leq-3$, and fixed initial $y_{2}$ values $y_{2,0}=-3.25$. An immediate observation is the dense complexity of the basins in three-dimensional space, certainly encapsulating the aforementioned seemingly random distribution of basin points without any obvious defined structure. To get a better view of this complex three-dimensional plot, we show eight select $(x_{1},x_{2})$ plane slices of the three-dimensional basins in Fig. 9. To get a sense of the level of detail in Figs. 8 and 9, a good comparison to make is between Figs. 6(b) and 9(b), which show the same two-dimensional basin slice. An observation from Fig. 8 is the relative abundance of black basin points around the $y_{1,0}=-3.25$ plane. Figs. 9(a)9(d) support this, and they also show that more interesting geometrical structures arise in this region as opposed to the random distributions in Figs. 9(e)9(h). This observation is to be expected since the region around the $y_{1,0}=-3.25$ plane corresponds to initial states with $y_{1,0}\approx y_{2,0}$, so it makes sense that initial states in this region will have a higher chance of synchronizing and being attracted to the spiking attractor immediately, especially considering that we are examining a small range of initial $x$ values. The more defined geometric structures in Figs. 9(a)9(d) also make sense from this perspective since the slow variables do not have to travel far to reach their eventual values, which we mentioned earlier is the reason for the apparent random scattering in Figs. 7(b) and 9(e)9(h).
Figure 8: A three-dimensional slice of the black (nonchaotic) basin of the asymmetrically coupled Rulkov neuron system with $-2\leq x_{1,0}\leq 2$, $-2\leq x_{2,0}\leq 2$, $-3.5\leq y_{1,0}\leq-3$, and fixed initial $y_{2}$ values $y_{2,0}=-3.25$
(a) $y_{1,0}=-3.24$
(b) $y_{1,0}=-3.25$
(c) $y_{1,0}=-3.26$
(d) $y_{1,0}=-3.27$
(e) $y_{1,0}=-3.15$
(f) $y_{1,0}=-3.2$
(g) $y_{1,0}=-3.3$
(h) $y_{1,0}=-3.35$
Figure 9: Select horizontal slices of Fig. 8
@@ V Fractal basin boundary analysis
In Figs. 6 and 7, it is clear that the basins of the nonchaotic spiking attractor and strange pseudo-attractor are not divided by a smooth boundary. This naturally leads us to suspect that the basin boundary $\Sigma$ between the white and black basins might be fractal, which would lead to uncertainty regarding which attractor a given initial state will end up in [27]. In this section, we use the uncertainty exponent method [46] to determine the dimension of the basin boundary $\Sigma$ and the degree to which final states (in the short term) are uncertain. Briefly, the uncertainty exponent method works as follows. Consider a dynamical system in $n$-dimensional state space, and let $\mathbf{x}_{0}$ be an initial state with a small uncertainty $\epsilon$, meaning the true initial state may be anywhere in the $n$-ball $|\mathbf{x}-\mathbf{x}_{0}|<\epsilon$. Let $A$ and $C$ be the only two attractors of the dynamical system, and let $\hat{A}$ and $\hat{C}$ be their associated basins of attraction. $\mathbf{x}_{0}$ will be attracted to either $A$ or $C$, but the uncertainty $\epsilon$ may cause the perturbed initial state to be in a different basin than $\mathbf{x}_{0}$. Then, we expect the probability that we will falsely predict where a random uncertain state in $|\mathbf{x}-\mathbf{x}_{0}|<\epsilon$ will end up, denoted by $\varrho(\epsilon)$, to follow a power law:
@@
where $\mathfrak{u}\leq 1$ is called the uncertainty exponent. Then, when $\mathfrak{u}$ is small, significantly minimizing initial state uncertainty will have only a small effect on final state uncertainty. It is shown in [46] that the uncertainty exponent $\mathfrak{u}$ is related to the dimension of the basin boundary dividing $\hat{A}$ and $\hat{C}$ by
@@
Returning to the asymmetrically coupled Rulkov neuron system, we present the results of our analysis of the basin boundary $\Sigma$ between the white and black basins using uncertainty exponents. Similar to how we classified the intersection of the neuron system’s basins with $S_{2}$ and $S_{4}$, we examine the uncertainty exponents and fractalization of both $\Sigma\cap S_{2}^{\prime}$ and $\Sigma\cap S_{4}^{\prime}$. We have already defined $S_{2}^{\prime}$ as a specific square-shaped subset of $S_{2}$ (Eq. (22)), and we will similarly define $S_{4}^{\prime}\subset S_{4}$ to be the four-dimensional hypercube
@@
to see if any of them end up in a different basin from $\mathbf{X}_{0}$.
@@
with an $R^{2}$ value of 0.986. This indicates that $\mathfrak{u}_{2}\approx 0.314$, and since this is less than one, final state uncertainty does indeed exist in $S_{2}^{\prime}$. By Eq. (39), the fractal dimension $d_{2}$ of $\Sigma\cap S_{2}^{\prime}$ is
@@
with an $R^{2}$ value of 0.967. This indicates $\mathfrak{u}_{4}\approx 0.037$, which is exceedingly small and indicates extreme final state sensitivity. To put this in perspective, the fact that $\varrho_{4}(\epsilon)$ is proportional to $\epsilon^{\mathfrak{u}_{4}}$ means that to reduce the uncertainty regarding which attractor an initial state in $S_{4}^{\prime}$ will be attracted to by a factor of 10, we will need to reduce the initial uncertainty $\epsilon$ by a factor on the order of $10^{27}$. By Eq. (39), the fractal dimension $d_{4}$ of the basin boundary $\Sigma\cap S_{4}^{\prime}$ is
$d_{4}=n-\mathfrak{u}_{4}\approx 4-0.037=3.963.$ (46)
This indicates that this basin boundary is extremely fractal; it has a comparable geometry to a true four-dimensional object even though it divides four-dimensional space. Even though Figs. 7(b) and 8 don’t show the entire four-dimensional set $S_{4}^{\prime}$, they justify this high fractal dimension, where it is clear that the boundary dividing the seemingly random distribution of white and black points must be close to the dimension of the system’s state space.
@@ VI Conclusions
The primary objective of this study was to analyze, classify, and quantify the geometries of the attractors, basins of attraction, and fractal basin boundaries of an asymmetrically electrically coupled system of two identical nonchaotic Rulkov neurons. We discovered the existence of a quasimultistability in the system, with a true nonchaotic spiking attractor and a chaotic spiking-bursting pseudo-attractor. Using box counting, we found that the chaotic pseudo-attractor was fractal but that its fractal dimension didn’t coincide with its Lyapunov dimension. We also classified the basins of attraction of the nonchaotic spiking attractor and the chaotic pseudo-attractor. We found that in a two-dimensional slice of the basins near the slow variable values of the attractors, the basin of the chaotic pseudo-attractor took up almost all of the state space slice, while the basin of the nonchaotic spiking attractor had a finite measure and occupied increasingly small fractions of state space far away from its attractor. However, in all of four-dimensional state space, we discovered that both basins occupied fixed fractions of it. Finally, analyzing the system’s uncertainty exponents and basin boundaries, we found extreme final state sensitivity requiring initial state uncertainty improvements on the order of $10^{27}$ to improve final state uncertainty by a factor of 10, as well as fractal basin boundaries with dimension $d\approx 3.96$ taking up almost all of four-dimensional state space.
@@
We believe that our findings of complex and interesting geometries in the basins of attraction of this asymmetrically coupled Rulkov neuron system can be applied to other neuron models and dynamical systems, and we hope to see further study done in the classification of basins of attraction and quantification of fractal basin boundaries and final state sensitivity in the context of continuous-time and discrete-time neuronal systems. We also suggest research be done in using the method of basin entropy [63] to analyze the final state uncertainty and basin geometries of neuron systems, which extends this work. Additionally, we suggest further research be done into the existence of quasimultistability and pseudo-attractors in dynamical systems, which are concepts we found to yield rich and complex dynamics and geometries in our study.
We conclude this paper with a discussion of the biological relevance of our results. Multistability is a prominent and important property in biological neuron systems that has been observed in a number of experimental studies [64, 65, 66]. One of the reasons why multistability is important for neuron function and living organisms is that its presence in cells can result in multipotentiality [67], which is an ability of certain cells to develop into multiple distinct cell types. This property can be modeled by a dynamical attraction, such as with cancer attractors [68, 69]. The multistability we observed emerging from the interaction of neurons in this study has been more broadly found to be involved in the pathogenesis of neurological diseases such as Parkinson’s disease [70, 71]. As one of the first studies on the emergence of final state sensitivity and complex basin geometry in neuronal systems, the results detailed in this paper could have important biological significance, with minute variations in the state of biological neurons leading to vastly different functional outcomes through mechanisms of multipotentiality or pathogenesis. Therefore, building on existing experimental studies [72, 73, 74], we encourage exploration in experimentally replicating our results regarding final state sensitivity and complex basin geometry with coupled biological neurons.
@@ Appendix A Jacobian Matrix Partition
Using this, we can compactify $J(\mathbf{X})$ into the form shown in Eq. (11).